OpenStudy (anonymous):

Subtract: (6x+1 )- (-7x+2)

OpenStudy (stormfire1):

Simplify and add the like terms: 6x + 7x = 13x 1-2 = -1 Answer will be 13x - 1

OpenStudy (anonymous):

(6x+1 )- (-7x+2) = 6x +7x +1 -2= 13x -1

OpenStudy (aroub):

same as that question! add like terms.. 6x+1+7x-2 (6x+7x) + (1-2) 13x -1

jhonyy9 (jhonyy9):

6x+1+7x-2=13x-1

OpenStudy (anonymous):

Thanks; what I am doing is double checking my work that I have already done and you all have been so helpfull

OpenStudy (anonymous):

Thanks; what I am doing is double checking my work that I have already done and you all have been so helpfull